# Analyzing Fairness of Computer Vision and Natural Language Processing Models

### v2

@@ 5 Results and Discussions
This section presents the results and outcomes of applying fairness libraries to computer vision (CV) and natural language processing (NLP) models. The primary goal of this analysis is to improve fairness metrics while maintaining or enhancing performance metrics. Detailed code and implementation specifics are provided in [33].
For the CV model, the sensitive features analyzed for potential bias are gender and ethnicity, with the target variable being the prediction of a person’s age group. For the NLP model, the target variable is the determination of whether a patient requires urgent intervention, based on a combination of the doctor’s findings, patient gender, and age group. The sensitive features in this case are gender and age group, with the aim of identifying and addressing any model bias related to these attributes.
The subsequent discussion highlights the results obtained from applying fairness libraries to both the CV and NLP models. Table 1 summarizes the strengths and weaknesses of the libraries we use in this analysis (Fairlearn and AIF360) in addition to another library called What-If-Tool created by Google which is not powerful enough to be used in this study.
Table 1: Pros and Cons of the three fairness libraries.
For Fairlearn, accuracy was used to evaluate the overall predictive performance of the income classification model, while demographic parity difference served as a fairness metric to measure disparities in outcomes across demographic groups. Mitigation strategies were implemented at three stages of the machine learning pipeline: preprocessing, in-processing, and postprocessing.
@@
The evaluation identified the best algorithms as those balancing high predictive accuracy with reduced bias, demonstrating the need for an integrated fairness approach to mitigate biases in structured datasets.
@@
5.1 Bias Mitigation in Computer Vision Models: A Comparative Analysis of Fairlearn and AIF360
@@ 5.1.1 Fairlearn Library
The Fairlearn library was employed to evaluate fairness in a computer vision model designed for age group classification. The dataset consisted of four age groups: 0–18, 18–30, 30–80, and above 80 years. Since Fairlearn does not support multi-class classification directly, metrics were computed separately for each age group. Results are summarized in Figure 1.
@@
Demographic Parity Difference (DPD): This measures the difference in selection rates among groups within a sensitive feature, such as ethnicity. A small DPD indicates reduced bias, implying the model treats all groups equally.
Equalized Odds Difference (EOD): This evaluates the quality-of-service harm by comparing the model’s performance (e.g., false positive rates and true positive rates) across groups within a sensitive feature.
@@
Results:
For the age group 30–80, a DPD of 55% was observed, indicating significant bias. In contrast, bias was minimal for groups above 80 years.
Mitigation algorithms were applied to reduce bias:
Correlation Remover: Reduced bias moderately with a minimal drop in accuracy from 66% to 63%.
Exponentiated Gradient: Further reduced bias but decreased accuracy more significantly.
Threshold Optimizer: Eliminated bias almost entirely but reduced accuracy to 18%, indicating a severe trade-off between fairness and performance.
Grid Search: Demonstrated minimal impact on bias and decreased accuracy to 45%.
@@ Results:
Insights:
Reducing bias often results in reduced model accuracy.
The Correlation Remover and Exponentiated Gradient algorithms provided a balance between bias reduction and performance retention, making them preferable for this study.
Figure 1: The results of applying Fairlearn to the computer vision model.
@@ 5.1.2 AIF360 Library
The AIF360 library was used to analyze bias in the same computer vision problem, leveraging its capability to handle multi-class classification directly. Results are summarized in Figure 2. Five fairness metrics were analyzed:
@@ Results:
The accuracy of the original model was 70%, with significant bias observed across metrics.
Mitigation algorithms were applied:
Reweighing and Disparate Impact Remover: Reduced bias effectively while maintaining the original accuracy.
Adversarial Debiasing: Also showed promising results with minimal accuracy degradation.
Post-Processing Algorithms (Equalized Odds Postprocessing and Reject Option Classification): Performed poorly, significantly increasing some fairness metrics and reducing accuracy to 17–18%.
@@
AIF360 provides a robust framework for bias analysis with multiple metrics and mitigation algorithms covering different stages of the machine learning lifecycle.
Reweighing and Disparate Impact Remover were the most effective algorithms, offering a balance between fairness and accuracy.
@@
This comparative analysis underscores the trade-offs between bias reduction and model accuracy in fairness-focused machine learning. While Fairlearn offers simplicity for binary tasks, AIF360 provides a more comprehensive suite for multi-class classification. The study highlights the importance of selecting appropriate mitigation algorithms based on the specific requirements and constraints of the application.
Figure 2: The results of applying AIF360 to the computer vision model.
5.2 Bias Mitigation in NLP Models: Evaluation Using Fairlearn and AIF360
@@ 5.2.1 Fairlearn Library
The Fairlearn library was utilized to analyze fairness in a binary classification task within a natural language processing (NLP) model. Two fairness metrics were evaluated: Demographic Parity Difference (DPD) and Equalized Odds Difference (EOD). Results are summarized in Figure 3.
@@ Results:
The original model showed a high performance with an accuracy of 97–98%.
Grid Search: Increased bias, making it the least effective mitigation algorithm.
Correlation Remover and Exponentiated Gradient: Showed minimal changes to the fairness metrics compared to the original model, indicating limited efficacy.
Threshold Optimizer: Demonstrated the best performance among the mitigation algorithms. It reduced the EOD while maintaining the original accuracy, though it did not affect the DPD.
@@
The Threshold Optimizer emerged as the most effective mitigation algorithm for reducing EOD without compromising performance.
Correlation Remover and Exponentiated Gradient had negligible impact on bias, while Grid Search was counterproductive, increasing bias metrics.
Figure 3: The results of applying Fairlearn to the NLP model.
@@ 5.2.2 AIF360 Library
The AIF360 library was similarly applied to assess fairness in the same NLP model, with the same metrics used for evaluation.
@@ Results:
The original model maintained an accuracy of 98%.
No observable changes in fairness metrics were noted after applying various mitigation algorithms.
The mitigation algorithms did not significantly influence either the fairness or performance metrics.
@@
The lack of improvement in fairness metrics suggests that the mitigation algorithms in AIF360 may not be impactful for this specific NLP model.
Alternative models or configurations may yield better results with the AIF360 library.
In the context of NLP models, the Threshold Optimizer from Fairlearn showed some potential in reducing bias while preserving accuracy, particularly for EOD. However, the overall impact of mitigation algorithms in both Fairlearn and AIF360 was limited, highlighting the need for further exploration of model-specific or task-specific fairness solutions.
Figure 4: The results of applying AIF360 to the NLP model.
@@ 6 Conclusion
The study highlighted the role of biased predictions in reinforcing systemic inequalities and underscores the need for evaluating and improving fairness in ML workflows. By utilizing a Kaggle dataset, this research provided practical insights into how fairness libraries, such as Fairlearn by Microsoft and AIF360 by IBM, can be employed to assess, visualize, and mitigate bias in ML models.
The findings indicate that each fairness tool has its own strengths and limitations in addressing bias, offering valuable guidance for researchers and practitioners. These insights contribute to the growing field of ML fairness, reinforcing the necessity of integrating fairness solutions into real-world ML applications to build more equitable and responsible systems. The study ultimately underscores the need for continuous efforts in ensuring that ML models operate ethically and do not perpetuate existing societal inequalities.

### v3

@@ 5 Results and Discussions
This section presents the results of applying fairness libraries to computer vision (CV) and natural language processing (NLP) models. The primary objective of this analysis is to improve fairness metrics while maintaining or enhancing performance metrics. Detailed code and implementation specifics are provided in [38].
For the CV model, the sensitive features analyzed for potential bias are gender and ethnicity, with the target variable being the prediction of a person’s age group. In the NLP model, the target variable is determining whether a patient requires urgent intervention, based on a combination of the doctor’s findings, patient gender, and age group. The sensitive features here are gender and age group, with the goal of identifying and addressing any model bias related to these attributes.
The following discussion highlights the results obtained from applying fairness libraries to both the CV and NLP models.
For Fairlearn, accuracy was used to evaluate the overall predictive performance, while demographic parity difference served as a fairness metric to measure disparities across demographic groups. Mitigation strategies were implemented at three stages of the ML pipeline: preprocessing, in-processing, and post-processing:
@@
The evaluation identified the most effective algorithms as those balancing high predictive accuracy with reduced bias, demonstrating the importance of an integrated fairness approach to mitigate biases in structured datasets.
@@
5.1 Bias Mitigation in Computer Vision Models
@@ 5.1.1 Fairlearn Library
The Fairlearn library was used to evaluate fairness in a computer vision model designed for age group classification. The dataset consisted of four age groups: 0–18, 18–30, 30–80, and above 80 years. Since Fairlearn does not support multi-class classification directly, metrics were computed separately for each age group. In Figure 2, we present the values of the demographic parity and equalized odds metrics for each subgroup of the sensitive feature (age group). The upper panel displays the demographic parity values for each subgroup, starting with the original model (accuracy = 0.66). We then show the metrics after applying different mitigation algorithms, such as Correlation Remover, Exponentiated Gradient, Grid Search, and Threshold Optimizer. The lower panel shows the equalized odds values for the same cases. Figure 3 presents the effect of applying the mitigation algorithms in sequential order. The results are discussed below.
@@
Demographic Parity Difference (DPD): This measures the difference in selection rates among groups within a sensitive feature, such as age group. A small DPD indicates reduced bias, meaning the model treats all groups equally.
Equalized Odds Difference (EOD): This evaluates the model’s performance, comparing false positive and true positive rates across groups within a sensitive feature.
@@
Original Model: Accuracy = 66%.
Demographic Parity Difference (DPD):
The highest DPD was 55%.
After applying the mitigation algorithms individually, the highest DPD decreased to 3.5% (best case scenario) with the Threshold Optimizer algorithm, but accuracy dropped to 18%.
When applying the mitigation algorithms sequentially, the highest DPD decreased to 5% (best case scenario) with the combination of Correlation Remover + Exponentiated Gradient, while accuracy increased to 30.1%.
Comparing the individual and sequential approaches: The individual application reduced bias by 93.63%, but accuracy dropped by 72.72%. The sequential application reduced bias by 90.91%, with a smaller accuracy decrease of 54.39%. This suggests that the sequential approach maintains model performance better than the individual approach while achieving similar levels of bias reduction.
Equalized Odds Difference (EOD):
The highest EOD was 31.1%.
After applying the mitigation algorithms individually, the highest EOD decreased to 3.4% (best case scenario) with the Threshold Optimizer algorithm, but accuracy dropped to 18%.
When applying the mitigation algorithms sequentially, the highest EOD decreased to 7.8% (best case scenario) with Correlation Remover + Exponentiated Gradient, and accuracy increased to 30.1%.
Comparing the individual and sequential approaches: The individual application reduced bias by 89.07%, but accuracy dropped by 72.72%. The sequential application reduced bias by 74.92%, with a smaller accuracy decrease of 54.39%. This shows that while the sequential approach maintains performance better, it does not reduce bias as effectively as the individual approach.
Metric | Accuracy of Original Model | Subgroup with Max Bias | Algorithm with Max Mitigation | Algorithm with Min Mitigation |
Demographic Parity Difference (DPD) (%) | 66% | 55% (30-80y) | Threshold Optimizer (, Acc = 18%) | Grid Search (, Acc = 45%) |
Equalized Odds Difference (EOD) (%) | 66% |
31.1%
(0-18 y) |
Threshold Optimizer (, Acc = 18%) | Correlation Remover (, Acc = 63%) |
Metric | Accuracy of Original Model | Subgroup with Max Bias | Algorithm with Max Mitigation | Algorithm with Min Mitigation |
Demographic Parity Difference (DPD) (%) | 66% | 55% (30-80y) | Correlation Remover + Exponentiated Gradient (, Acc = 30.10%) | Correlation Remover + Grid Search (, Acc = 51.70%) |
Equalized Odds Difference (EOD) (%) | 66% |
31.1%
(0-18 y) |
Correlation Remover + Exponentiated Gradient (, Acc = 30.1%) | Correlation Remover + Grid Search (, Acc = 51.7%) |
@@ 5.1.2 AIF360 Library
The AIF360 library is used to analyze bias in the same computer vision problem, taking advantage of its ability to handle multi-class classification directly. In Figure 4, we present various fairness metrics from the AIF360 library, including Statistical Parity Difference, Average Odds Difference, Equal Opportunity Difference, Theil Index, and Generalized Entropy Index. First, we show the values of these metrics for the original model, which has an accuracy of 0.7050, without applying any mitigation algorithms. Then, we demonstrate the impact of applying different mitigation algorithms on both fairness metrics and model accuracy. The algorithms used include Reweighing, Disparate Impact Remover, Learning Fair Representations, Adversarial Debiasing, Equalized Odds Postprocessing, and Reject Option Classification. In Figure 5, we show the effect of applying these mitigation algorithms in sequential order, highlighting their influence on both model bias and performance. The results are discussed below. Five fairness metrics are analyzed:
@@ Results:
The results of the AIF360 metrics with applying the mitigation algorithms one at a time (individually) and in sequential order are shown on Figure 4,5 and the summary of the results are shown in Table (6). -
@@
From the results in Table (6), we find that the original model has accuracy of 70.5% and the highest bias metric (Theil Index) is 58.61%.After applying the mitigation algorithms individually, the highest bias metric (Theil Index) decreased to 47.14% in the best-case scenario, with accuracy reaching 70.90% using the Disparate Impact Remover algorithm. In contrast, when applying the mitigation algorithms sequentially, the highest bias metric (Theil Index) dropped to 53.30%, with accuracy at 62.83% using the combination of Reweighing + Disparate Impact Remover + Adversarial Debiasing + Reject Option Classification (RW+DIR+AD+ROC). Comparing the two approaches, the individual application reduced bias by 19.57% while maintaining nearly the same accuracy. Meanwhile, the sequential application reduced bias by 9.06%, but accuracy decreased by 10.88%. These results indicate that the individual application outperforms the sequential approach in bias mitigation.
@@
This comparative analysis underscores the trade-offs between bias reduction and model accuracy in fairness-focused ML. While Fairlearn offers simplicity for binary tasks, AIF360 provides a more comprehensive suite for multi-class classification. The study highlights the importance of selecting appropriate method of applying the mitigation algorithms to obtain the highest efficiency by reducing the ML bias and maintaining the model performance.
Mitigation Algorithm | Accuracy of Original Model | Metric with Max Bias | Algorithm with Max Mitigation | Algorithm with Min Mitigation |
Individual Application | 70.5% | 58.61% (Theil Index) | Disparate Impact Remover (Theil Index=47.14%, Acc = 70.90%) | Equalized Odds Postprocessing (Theil Index = 63.12%, Acc = 17.86%) |
sequential Application | 70.5% | 58.61% (Theil Index) | RW+DIR+AD+ROC (Theil Index = 53.30%, Acc = 62.83%) | RW+DIR+AD+EOP (Theil Index = 54.56%, Acc = 66.90%) |
5.2 Bias Mitigation in NLP Models
@@ 5.2.1 Fairlearn Library
The Fairlearn library was used to assess fairness in a binary classification task within a natural language processing (NLP) model. Two fairness metrics were evaluated: Demographic Parity Difference (DPD) and Equalized Odds Difference (EOD). Since the model is binary, we present the metrics for each class. In Figure 6, we show the values of these metrics for the original model, which has an accuracy of 0.981. We then demonstrate the effects of applying various mitigation algorithms, including Correlation Remover, Exponentiated Gradient, Grid Search, and Threshold Optimizer, on both fairness and performance metrics. Figure 7 presents the results of applying the mitigation algorithms sequentially. The discussion of the results can be found below.
@@ Results:
The results of the Fairlearn metrics with applying the mitigation algorithms one at a time (individually) and in sequential order are shown on Figure 6,7 and the summary of the results are shown in Table (7). -
@@
From the results in Table (7), we observe that the original model has an accuracy of 98.1% and the highest bias metric (DPD) is 27.4%.After applying the mitigation algorithms individually, the highest bias metric (DPD) decreased to 21.4% with the Exponentiated Gradient algorithm, while the accuracy dropped to 95.4%. On the other hand, when the mitigation algorithms were applied in a sequential order, the highest bias metric (DPD) decreased to 17.9%, with an accuracy of 92.7% using the combination of Correlation Remover and Exponentiated Gradient.
Comparing the individual and sequential applications of the algorithms, we find that the individual application reduces bias by 21.89%, but the model’s accuracy decreases by 2.75%. In contrast, the sequential application reduces bias by 34.67%, but the model’s accuracy decreases by 5.5%. These results show that the sequential application is more effective in reducing bias, while maintaining model performance at a similar level.
Mitigation Algorithm | Accuracy of Original Model | Metric with Max Bias | Algorithm with Max Mitigation | Algorithm with Min Mitigation |
Individual Application | 98.1% | 27.4% (DPD) | Exponentiated Gradient (DPD = 21.4%, Acc = 95.4%) | Grid Search (EOD = 58.3%, Acc = 90%) |
sequential Application | 98.1% | 27.4% (DPD) | Correlation Remover + Exponentiated Gradient (DPD = 17.9%, Acc = 92.7%) | Correlation Remover + Grid Search (EOD = 41.3%, Acc = 93.6%) |
@@ 5.2.2 AIF360 Library
The AIF360 library was also applied to evaluate fairness in the same NLP model, using the same fairness metrics for assessment. Figure 8 presents various fairness metrics available in the AIF360 library, such as statistical parity difference, average odds difference, equal opportunity difference, Theil index, and generalized entropy index. Initially, we display the values of these metrics for the original model, which achieves an accuracy of 98.14%, without applying any mitigation algorithms. Following this, we analyze the impact of different mitigation algorithms on both fairness metrics and model accuracy. The mitigation techniques used include reweighing, disparate impact remover, learning fair representations, adversarial debiasing, equalized odds postprocessing, and reject option classification. Figure 9 shows the effect of applying these mitigation algorithms in a sequential order. The summarized results, highlighting the analysis of these five fairness metrics, are shown in Table 8.
@@ Results:
The results of the AIF360 metrics with applying the mitigation algorithms one at a time (individually) and in sequential order are shown on Figure 8,9 and the summary of the results are shown in Table (8). -
@@
From the results in Table (8), we find that the original model has accuracy of 98.14% and the highest bias metric (SPD) is 19.80%.After applying the mitigation algorithms individually, the highest bias metric (SPD) decreased to 19.71%, with the model’s accuracy reaching 97.23% using the Reweighing algorithm. In contrast, when the mitigation algorithms were applied in a sequential order, the highest bias metric (SPD) remained unchanged, while the accuracy slightly dropped to 97.32% with the combination of Reweighing + Disparate Impact Remover + Adversarial Debiasing + Equalized Odds Postprocessing (RW+DIR+AD+EOP). A comparison of the individual and sequential application of algorithms reveals that both methods resulted in minimal change to the model’s bias or performance.
Mitigation Algorithm | Accuracy of Original Model | Metric with Max Bias | Algorithm with Max Mitigation | Algorithm with Min Mitigation |
Individual Application | 98.14% | 19.79% (SPD) | Reweighing (SPD = 19.71%, Acc = 97.23%) | Learning Fair Representations (SPD = 25.81%, Acc = 72.89%) |
sequential Application | 98.14% | 19.80% (SPD) | RW+DIR+AD+EOP (SPD = 19.80%, Acc = 97.32%) | RW+DIR+AD+ROC (SPD = 20.16%, Acc = 97.46%) |
6 Future Work
One of the key challenges in applying fairness interventions to generative AI models is the complexity of their architectures and outputs [42]. Unlike traditional models, generative models produce unstructured data such as text, images, and other media, making it more difficult to detect and mitigate bias [43]. To address this, fairness metrics specifically designed for generative models—such as distribution-level evaluations and fairness-aware assessments of generation quality—need to be developed. Additionally, the multi-modal nature of these models requires advancements in fairness algorithms to handle diverse data types effectively. A major obstacle remains the lack of benchmark datasets explicitly designed for fairness evaluation in generative AI. Creating such datasets and integrating them into fairness libraries would enable more robust testing and comparison of mitigation strategies.
The techniques discussed in this paper can be extended to generative AI models, including Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and transformer-based models like GPT and DALL-E. Applying sequential mitigation algorithms across the three stages of the ML lifecycle would help identify and reduce biases in both their training processes and outputs. This approach aligns with recent findings that mitigation strategies can be effective when applied holistically across the ML lifecycle [44]. It holds significant potential in areas like content creation, data augmentation, and conversational AI, where fairness and ethical considerations are critical [45]. Another important extension is exploring how sequential fairness interventions impact the creative capacities of generative models, ensuring diversity and representativeness are preserved without compromising quality [46].
Extending fairness interventions to generative AI models has a broader impact on promoting equity and inclusivity in AI-generated content. Generative models are increasingly used in domains such as art, education, entertainment, and media, where biased outputs could perpetuate harmful stereotypes or exclusion [47]. Addressing fairness in these models ensures that they contribute to a more inclusive digital ecosystem [48]. This research will provide actionable insights and tools for practitioners to incorporate fairness into generative AI systems, fostering trust and accountability in AI deployment [48]. Ultimately, applying fairness principles to generative models will transform how these systems are designed, evaluated, and utilized, ensuring that their outputs align with ethical standards and societal values.
Explainability in generative AI systems remains a significant challenge, making it difficult to understand how biases are introduced and propagated. To address this, integrating Explainable AI (XAI) techniques is essential. These techniques can reveal the inner workings of generative models and enable targeted fairness interventions at key points [49]. Furthermore, the absence of benchmark datasets for evaluating fairness in multi-modal generative models underscores the need for diverse, high-quality datasets that capture potential biases and facilitate rigorous testing and validation of fairness strategies [43]. XAI can also enhance model clarity, offering stakeholders insights into how fairness interventions influence outcomes [50]. Future research could explore fairness-aware generation in multi-modal conversational agents, ensuring equitable interactions across languages, cultures, and user demographics [46]. By integrating XAI into generative models, users can understand and challenge biases, promoting accountability and ethical AI practices. Extending fairness interventions to multi-modal and generative systems supports global efforts to develop AI that reflects societal values and ethical standards, contributing to a more equitable digital ecosystem.
@@ 7 Conclusion
In this work, we studied the fairness of the machine learning models with using unstructured data. We used the fairness libraries Fairlearn by Microsoft and AIF360 by IBM, and used Kaggle datasets including images with a computer vision model and text with natural language processing model. This research provided practical insights of a comparative study between applying the mitigation algorithms one at a time (individually) in the stages of the ML lifecycle and in more than a stage in a sequential order in both models using both libraries. The study highlighted the role of biased predictions in reinforcing systemic inequalities and underscores the need for evaluating and improving fairness in ML workflows.
From the results of applying the fairlearn and AIF360 mitigation algorithms to the computer vision model, we found that fairlearn reduced the bias in the model substantially better than AIF360, but at the cost of dropping the model performance. By applying the mitigation algorithms in a sequential order, fairlearn was able to maintain the model performance by about 67.22% over that of applying the algorithms one at a time with reducing the bias almost to the same level. On the other hand, the individual application of the algorithms in AIF360 library showed better performance over the sequential order application on the level of the model fairness and performance.
In the case of the NLP model, we found that fairlearn reduced the bias in the model better than AIF360 which nearly did not change neither the fairness nor the performance of the model. By applying the mitigation algorithms in a sequential order, fairlearn was able to slightly reduce the bias better than applying individually by a factor of 16.35%.
Our findings indicated that applying mitigation algorithms in a sequential order can produce favorable results compared to using a single algorithm. This approach successfully reduced bias while preserving model performance in some cases, as reflected by minimal changes in accuracy. By showing that a sequential strategy can balance fairness and performance. Our work introduced a novel and practical framework for research and industrial efforts in mitigating ML bias.